A certain Ruby laser emits 100J pulses of light whose wavelength is 6940A.What is the minimum no. of Cr3+ ions in the laser?
1
Expert's answer
2020-07-06T15:40:26-0400
"n=\frac{E}{E_\gamma}\\
\\E=\frac{hc}{\lambda}=\frac{1.986\cdot10^{-25}}{(694\cdot10^{-9})}=2.86\cdot10^{-19}J\\n=\frac{100}{2.86\cdot10^{-19}}=3.5\cdot10^{20}"